### v1

@@ 6.3.2 Audio Deepfake
Audio-related benchmark detection tasks are primarily concerned with assessing whether an audio segment belongs to Deepfake content. Such benchmarks typically encompass an extensive collection of Deepfake audio produced by technologies like TTS and VC. Detection methods must be capable of discerning the discrepancies between these and authentic audio content, or determining through specific characteristics whether the audio has been synthesized using Deepfake or analogous technologies.
@@
EmoFake [302]. This benchmark innovatively proposes the generation of fake audio by modifying the emotional state of the source audio. The construction process involves sampling data from the ESD dataset and utilizing open-source Emotional Voice Conversion (EVC) models, such as VAW-GAN-CWT and DeepEST, to transform the original emotion into another. The result is a dataset comprising 39,900 fake data samples and 17,500 real data samples. For the evaluation process, this benchmark employs the Equal Error Rate (ERR) metric to assess the performance of detection methods.
@@
Wavefake [64]. This benchmark’s construction process starts by sampling data from two datasets, LJSPEECH and JSUT, and uses six state-of-the-art architectures across two languages to generate corresponding recreated fake data. Ultimately, this yields a dataset containing 177,985 generated audio samples. For the evaluation process, the benchmark uses the ERR metric to evaluate the performance of detection methods.
@@
LibriSeVoc [227]. This benchmark focuses on vocoder artifact detection tasks. It is built upon the LibriTTS speech corpus, utilizing six SOTA neural vocoders to generate samples, each synthesized sample will contain specific vocoder artifacts. The resulting dataset includes 13,201 real audio samples and 79,206 fake audio samples. For the evaluation process, the benchmark uses the ERR metric to evaluate the performance of detection methods.
@@
Real-and-fake [198]. For this benchmark, the construction process starts with a specific phrase dataset and employs multiple TTS systems to generate synthetic utterances as fake samples. It also collects real human utterances from various open-source datasets and internet recordings to serve as true samples. The final dataset contains more than 87,000 fake samples and more than 111,000 real samples. Depending on the preprocessing method, the dataset can be divided into four subsets: for-rece, for-2-sec, for-norm, and for-original. The evaluation process uses both ERR and Accuracy metrics to assess the performance of detection methods.
@@
In-the-Wild [166]. This benchmark emphasizes evaluating the generalization capabilities of related work in real-world scenarios. It is constructed by collecting a large amount of audio data from public channels, involving celebrities and politicians. Fake samples are clipped from publicly available Deepfake files, while real samples are gathered by collecting genuine materials such as speeches and podcasts from the same speakers featured in the fake samples. The final dataset includes 19,963 fake samples and 11,816 real samples, For the evaluation process, the benchmark utilizes the ERR metric to evaluate the performance of detection methods.
@@
ADD2023 [276]. This benchmark comprises three subsets: FG-D, LR, and AR. FG-D focuses on the identification of fake utterances, LR on locating the manipulated regions in partially fake audio, and AR aims to recognize the algorithms used in Deepfake utterances. FG-D contains 172,819 real samples and 113,042 fake samples totaling 285,861 samples; LR has 46,554 real samples and 65,449 fake samples totaling 112,003 samples; AR includes 14,907 real samples and 95,383 fake samples totaling 110,290 samples. The evaluation process utilizes F1-score, EER, and A-sentence metrics to assess the performance of detection methods.
@@
ASVspoof 2021 [266]. This benchmark is based on the VCTK base corpus and some undisclosed corpora, divided into three subsets: LA, PA, and DF, with the DF subset being relevant to Deepfake detection. The DF subset focuses on the detection of manipulated, compressed speech data posted online, reflecting a scenario where an attacker has access to the voice data of a targeted victim. The DF subset includes 22,617 real samples and 589,212 fake samples. For the evaluation process, the benchmark employs the ERR metric to evaluate the performance of detection methods.
@@ 7 Challenges and Future Directions
Multimodal Deceptive FAIGC Detect. With the rapid evolution of AIGC technologies, a new challenge emerges: malicious users increasingly exploit AI to craft highly convincing fake multimodal content, thereby significantly complicating public efforts in discerning and counteracting misinformation. Recent research primarily addresses the detection of fake content generated by LLMs in textual form [93, 250, 41, 73, 179], yet this scrutiny has not been sufficiently extended to multimodal domains. By its complex and realistic nature, Multimodal fake content presents a more daunting detection challenge than its textual counterpart. While some studies have examined the outputs of generative models from a multimodal standpoint [137], there remains a shortfall in investigating the veracity of these outputs. The multimodal fake content detection field is notably lacking in comprehensive research and datasets, a gap that warrants significant academic attention and exploration. Concurrently, while deceptive content detection has predominantly focused on fake news, lesser attention has been given to other prevalent forms such as rumors, clickbait, and cherry-picking. Given the ubiquity of these types of content, developing targeted detection algorithms and establishing evaluative datasets represent another critical research avenue.

### v2

@@ 6.3.2 Audio Deepfake
Audio-related benchmark detection tasks primarily assess whether an audio segment belongs to Deepfake content. Such benchmarks typically encompass an extensive collection of Deepfake audio produced by technologies like TTS and VC. Detection methods must be capable of discerning the discrepancies between these and authentic audio content, or determining through specific characteristics whether the audio has been synthesized using Deepfake or analogous technologies.
@@
EmoFake [302]. This benchmark innovatively proposes the generation of fake audio by modifying the emotional state of the source audio. The construction process involves sampling data from the ESD dataset and utilizing open-source Emotional Voice Conversion (EVC) models, such as VAW-GAN-CWT and DeepEST, to transform the original emotion into another. The result is a dataset comprising 39,900 fake data samples and 17,500 real data samples.
@@
Wavefake [64]. This benchmark’s construction process starts by sampling data from two datasets, LJSPEECH and JSUT, and uses six state-of-the-art architectures across two languages to generate corresponding recreated fake data. Ultimately, this yields a dataset containing 177,985 generated audio samples.
@@
LibriSeVoc [227]. This benchmark focuses on vocoder artifact detection tasks. It is built upon the LibriTTS speech corpus, utilizing six SOTA neural vocoders to generate samples, each synthesized sample will contain specific vocoder artifacts. The resulting dataset includes 13,201 real audio samples and 79,206 fake audio samples.
@@
Real-and-fake [198]. For this benchmark, the construction process starts with a specific phrase dataset and employs multiple TTS systems to generate synthetic utterances as fake samples. It also collects real human utterances from various open-source datasets and internet recordings to serve as true samples. The final dataset contains more than 87,000 fake samples and more than 111,000 real samples. Depending on the preprocessing method, the dataset can be divided into four subsets: for-rece, for-2-sec, for-norm, and for-original.
@@
In-the-Wild [166]. This benchmark emphasizes evaluating the generalization capabilities of related work in real-world scenarios. It is constructed by collecting a large amount of audio data from public channels, involving celebrities and politicians. Fake samples are clipped from publicly available Deepfake files, while real samples are gathered by collecting genuine materials such as speeches and podcasts from the same speakers featured in the fake samples. The final dataset includes 19,963 fake samples and 11,816 real samples.
@@
ADD2023 [276]. This benchmark comprises three subsets: FG-D, LR, and AR. FG-D focuses on the identification of fake utterances, LR on locating the manipulated regions in partially fake audio, and AR aims to recognize the algorithms used in Deepfake utterances. FG-D contains 172,819 real samples and 113,042 fake samples; LR has 46,554 real samples and 65,449 fake samples; AR includes 14,907 real samples and 95,383 fake samples.
@@
ASVspoof 2021 [266]. This benchmark is based on the VCTK base corpus and some undisclosed corpora, divided into three subsets: LA, PA, and DF, with the DF subset being relevant to Deepfake detection. The DF subset focuses on the detection of manipulated, compressed speech data posted online, reflecting a scenario where an attacker has access to the voice data of a targeted victim. The DF subset includes 22,617 real samples and 589,212 fake samples.
@@ 7 Challenges and Future Directions
Multimodal Deceptive FAIGC Detect. With the rapid evolution of AIGC technologies, a new challenge emerges: malicious users increasingly exploit AI to craft highly convincing fake multimodal content, thereby significantly complicating public efforts in discerning and counteracting misinformation. Recent research primarily addresses the detection of fake content generated by LLMs in textual form [93, 250, 41, 73, 179], yet this scrutiny has not been sufficiently extended to multimodal domains. By its complex and realistic nature, Multimodal fake content presents a more daunting detection challenge than its textual counterpart. While some studies have examined the outputs of generative models from a multimodal standpoint [137], there remains a shortfall in investigating the veracity of these outputs. The multimodal fake content detection field lacks comprehensive research and datasets, which warrants significant academic attention and exploration.